A b and c share profit on a business deal a receive 1/3 of the profit b receive 2/3 of the remainder if Ope receive 12000 how much profit did they share
Found 2 solutions by saw, ikleyn:
Answer by saw(34)

a = 1/3*12000
b = 2/3*(12000-a)
c =12000-(a+b)
a+b+c = 12000
Use substitution
4000+5333.333+2666.667 = 12000
So A received 4000, B received 5333.333, and C received 2666.667.

Answer by ikleyn(52848)

.
A b and c share profit on a business deal a receive 1/3 of the profit b receive 2/3 of the remainder
if Ope receive 12000 how much profit did they share
~~~~~~~~~~~~~~~~~~~


1)   Who is Ope ?

2)   As worded,  printed,  submitted and presented,  this post  MAKES  no  SENSE.


It is not a  Math problem - - - it is a  FAKE,  instead.

------------

IGNORE the "solution" by @saw,
since it is as NONSENSICAL, as your post is.


/\/\/\/\/\/\/

Yesterday a  HUGE  amount of nonsensical posts was submitted to the forum.

I decided to make a list of nonsensical posts that come to the forum.


From time to time,  I will send this list to the managers of this project,
asking them to take  MEASURES  against the authors of such posts.

The given post will be  #2  in this list.